40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 65 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  14375   Thu Dec 20 21:29:41 2018 JonOmnistructureUpgradeVacuum Controls Switchover Completed

[Jon, Chub, Koji, Gautam]

Summary

Today we carried out the first pumpdown with the new vacuum controls system in place. It performed well. The only problem encountered was with software interlocks spuriously closing valves as the Pirani gauges crossed 1E-4 torr. At that point their readback changes from a number to "L OE-04, " which the system interpreted as a gauge failure instead of "<1E-4." This posed no danger and was fixed on the spot. The main volume was pumped to ~10 torr using roughing pumps 1 and 3. We were limited only by time, as we didn't get started pumping the main volume until after 1pm. The three turbo pumps were also run and tested in parallel, but were isolated to the pumpspool volume. At the end of the day, we closed every pneumatic valve and shut down all five pumps. The main volume is sealed off at ~10 torr, and the pumpspool volume is at ~1e-6 torr. We are leaving the system parked in this state for the holidays. 

Main Volume Pumpdown Procedure

In pumping down the main volume, we carried out the following procedure.

  1. Initially: All valves closed (including manual valves RV1 and VV1); all pumps OFF.
  2. Manually connected roughing pump line to pumpspool via KF joint.
  3. Turned ON RP1 and RP2.
  4. Waited until roughing pump line pressure (PRP) < 0.5 torr.
  5. Opened V3.
  6. Waited until roughing pump line pressure (PRP) < 0.5 torr.
  7. Manually opened RV1 throttling valve to main volume until pumpdown rate reached ~3 torr/min (~3 hours on roughing pumps).
  8. Waited until main volume pressure (P1a/P1b) < 0.5 torr.

We didn't quite reach the end of step 8 by the time we had to stop. The next step would be to valve out the roughing pumps and to valve in the turbo pumps.

Hardware & Channel Assignments

All of the new hardware is now permanently installed in the vacuum rack. This includes the SuperMicro rack server (c1vac), the IOLAN serial device server, a vacuum subnet switch, and the Acromag chassis. Every valve/pump signal cable that formerly connected to the VME bus through terminal blocks has been refitted with a D-sub connector and screwed directly onto feedthroughs on the Acromag chassis.

The attached pdf contains the master list of assigned Acromag channels and their wiring.

Attachment 1: 40m_vacuum_acromag_channels.pdf
40m_vacuum_acromag_channels.pdf 40m_vacuum_acromag_channels.pdf 40m_vacuum_acromag_channels.pdf
  14374   Thu Dec 20 17:17:41 2018 gautamUpdateCDSLogging of new Vacuum channels

Added the following channels to C0EDCU.ini:

[C1:Vac-P1b_pressure]
units=torr
[C1:Vac-PRP_pressure]
units=torr
[C1:Vac-PTP2_pressure]
units=torr
[C1:Vac-PTP3_pressure]
units=torr
[C1:Vac-TP2_rot]
units=kRPM
[C1:Vac-TP3_rot]
units=kRPM

Also modified the old P1 channel to

[C1:Vac-P1a_pressure]
units=torr

Unfortunately, we realized too late that we don't have these channels in the frames, so we don't have the data from this test pumpdown logged, but we will have future stuff. I say we should also log diagnostics from the pumps, such as temperature, current etc. After making the changes, I restarted the daqd processes.


Things to add to ASA wiki page once the wiki comes back online:

  1. What is the safe way to clean the cryo pump if we want to use it again?
  2. What are safe conditions to turn the RGA on?
  14373   Thu Dec 20 10:28:43 2018 gautamUpdateVACHeavy doors back on for pumpdown 82

[Chub, Koji, Gautam]

We replaced the EY and IOO chamber heavy doors by 10:10 am PST. Torquing was done first oen round at 25 ft-lb, next at 45 ft-lb (we trust the calibration on the torque wrench, but how reliable is this? And how important are these numbers in ensuring a smooth pumpdown?). All went smooth. The interior of the IOO chamber was found to be dirty when Koji ran a wipe along some surfaces.

For this pumpdown, we aren't so concerned with having the IFO in an operating state as we will certainly vent it again early next year. So we didn't follow the full close-up checklist.

Jon and Chub and Koji are working on starting the pumpdown now... In order to not have to wear laser safety goggles while we closed doors and pumped down, I turned off all the 1064nm lasers in the lab.

  14372   Thu Dec 20 08:38:27 2018 JonUpdateVACPumpdown tomorrow

Linked is the pumpdown procedure, contained in the old 40m documentation. The relevant procedure is "All Off --> Vacuum Normal" on page 11.

Quote:

I just spoke to Jon who asked me to make this elog - we will be ready to test one or more parts of the pumpdown procedure tomorrow (12/20), so we should proceed as planned to put the heavy doors back on EY and OMC chambers at 9am tomorrow morning. Jon will circulate a more detailed procedure about the pumpdown steps later today evening.

 

  14371   Wed Dec 19 22:11:28 2018 KojiUpdateGeneralHow to align the copper OMC

The OMC input optics layout is attached

Checked the spot position on OMMT-FM1. It was off from the center. This was causing the spot on OMMT1 off-center. This was fixed by the steering mirror for the AUX laser.

The beam alignment onto the OMC was tweaked with OMC-SM1 and OMC-SM2. This was the painful part. We had to make a sensor card that could get in to the narrow space of the OMC. (Attachment 2 right)

Attachment 2 left shows the naming convention of the OMC mirrors.

For the alignment, we gave 5Vpp trig waves at 3.1Hz to the input of the PZT amp so that the cavity is kept scanned continuously. Firstly check the rough spot positions for OMC-CM1 and OMC-CM2. If you carefully use the card, you can check if the beam is returning to OMC-IC. This return beam should have roughtly same hight as the incident beam. This can be adjusted by either of the steering mirrors.

Once the beam is going around the mirrors multiple times, the spot alignment can be checked at OMC-CM1. Bring a card right in front of CM1. If the card is lifter slightly above the incident spot, this automatically allows for the outgoing beam to go through. Depending on the pitch alignment, the next roundtrip (1RT) will be seen on the card. As you lift the card up more, you will be able to see more round trip beams (e.g. 2RT, 3RT, in the figure). If the yaw alignment is perfect, these spots would be lined up vertically. So you can try to align the horizontal direction with the steering mirrors. Then the vertical alignment can be done with the pitch knobs.

At this point you should be able to see some super high-order transmission at the OMC trans. For today, we stopped here as we already ran out of the knob ranges at multiple knobs. This is because the beam height in the mode matching telescope was not right, and the steering mirrors had to work more than their range.

Attachment 1: 110804_40m_OMC_layout.pdf
110804_40m_OMC_layout.pdf
Attachment 2: OMC_alignment.pdf
OMC_alignment.pdf
  14370   Wed Dec 19 21:14:50 2018 gautamUpdateVACPumpdown tomorrow

I just spoke to Jon who asked me to make this elog - we will be ready to test one or more parts of the pumpdown procedure tomorrow (12/20), so we should proceed as planned to put the heavy doors back on EY and OMC chambers at 9am tomorrow morning. Jon will circulate a more detailed procedure about the pumpdown steps later today evening.

  14369   Wed Dec 19 19:51:19 2018 gautamUpdateGeneralPumpdown prep

[Koji, gautam]

Summary:

We are ready to put the heavy doors back on the chambers and do some test pumpdowns tomorrow morning if Jon gives us the go-ahead. Also, Koji made the OMC resonate some of the AUX beam light we send into ityes

Details:

  1. EY work:
    • IMC was locked, and we attempted to locate the beam with an IR card inside the chamber.
    • Koji found that the beam was too high, we were over-shooting the entire black-glass baffle on the EY table.
    • So I moved the TTs to try and center the beam through the aperture of aforementioned baffle.
    • Once this was done, we found that the beam was misaligned in yaw by ~1-inch in transmission on the EY optics table (there was an iris in place marking the cavity transmission axis). This explains why I couldn't find any TRY flashes while moving the TTs around.
    • We hypothesize that without the 2 degree ETM wedge in place, there isn't a compatible axis for the ITM transmission to also make it through the EY baffle and transmission iris. Over ~1m, the 2 degree wedge makes roughly 1.4 inch translation in yaw, so this seems to be a plausible hypothesis.
    • The ETMY suspension was moved from the mini-cleanroom setup back into the EY vacuum chamber. Two clamps (finger tightened only) hold it in place on the NE edge of the optical table. We decided that this is a better resting palce for the cage over the holidays than an in-air cleanroom.
  2. OMC chamber work:
    • While we were in clean garb, we decided to also investigate the OMC situation a bit.
    • It quickly became apparent that it was hopeless for me to work in chamber in the tightly confined IOO chamber. So Koji went in to have a look.
    • Koji will post the detailed alignment procedure - but after some alignment of the AUX laser input beam axis using in air steering mirrors and Koji's expert tweaking of the pointing into the OMC, we observed some resonances of the OMC.
    • Attachment #1 shows the full-range triangle ramp applied to the OMC length PZT (top row) and the OMC REFL signal (bottom row), measured using a PDA520 (chosen for its large active area) connected to a scope (AC-coupled, 1Mohm impedance, averaged to make the dips more prominent).
    • The OMC transmission was also (barely) visible on an IR card.
    • So the OMC length PZT seems capable of sweeping the length of the cavity. Based on the size of the dips we saw, the MM into the cavity is sub 1-percent.
    • The transmission PDs didn't output any measurable signal - but I'm not sure that the satellite box / readout electronics have been carefully characterized on the electroncis bench, so that will have to be done first.
    • We replaced the copper cover of the OMC (finger tightened for now) in case we do any test pumpdowns tomorrow. HV supply has been turned off, and the AUX laser has been reverted to standby mode.
Attachment 1: OMCscan.pdf
OMCscan.pdf
  14368   Wed Dec 19 15:15:56 2018 gautamUpdateIOOTT1/TT2 stepping

I removed the ND filter from the ETMYT camera to facilitate searching for a TRY beam. This should be replaced before we go back to high power.

  14367   Wed Dec 19 14:19:15 2018 KojiSummaryVACPlan for pumpoing down test

We still need elaborated test procedure posted

12/29 Wed

  • Jon continues to work on valve actuator tests.
  • Chub continues to work on wiring / fixing wiring.
  • At the end of the day Jon is going to send out a notification email of "GO"/"NO GO" for pumping.

 

12/30 Thu

  • 9AM: Start closing two doors unless Jon gives us NO GO sign.
  • 10AM: Start pumping down
    • Test roughing pump capability via new control system
    • (Independently) Test turbo rotating procedure. This time we will not open the gate valve between the TP1 and the main volume. This is because we want to take care of the backing turbo loads while we gradually open the gate valve. This will take more hours to be done and we will not be able to finish this test by the end of Thu.
    • At the end of the procedure, we isolate the main volume, stop all the pumps, and vent the roghing pumps to save them from the oil backstream.

gautam: Koji and I were just staring at the vacuum screen, and realized that the drypumps, which are the backing pumps for TP2 and TP3, are not reflected on the MEDM screen. This should be rectified.

Steve also mentioned that the new small turbo controller does not directly interface with the drypump. So we need some system to delay the starting of the turbo itself, once the drypump has been engaged. Does this system exist?

Attachment 1: Screenshot_from_2018-12-19_14-49-34.png
Screenshot_from_2018-12-19_14-49-34.png
  14366   Wed Dec 19 00:12:46 2018 gautamUpdateOMC40m OMC DCC node

I made a node to collect drawings/schematics for the 40m OMC, added the length drive for now. We should collect other stuff (TT drivers, AA/AI, mechanical drawings etc) there as well for easy reference.

Some numbers FTR:

  • OMC length PZT capacitance was measured to be 209 nF.
  • Series resistance in HV path of OMC lenght PZT driver is 10 kohms, so this creates a LP with corner 1/2/pi/10kohm/200nF ~80 Hz.
  • Per Rob's thesis, the length PZT has DC actuation coefficient of 8.3 nm/V, ∼ 2 µm range. 
  14365   Tue Dec 18 18:13:32 2018 aaronUpdateOMCOMC L HV Piezo driver tests (again)

I tested the OMC-L HV driver box again, and made the following observations:

  • Drove the HV diff pins (2,7) with a 5V triangle wave
    • Observed that with a ~0.4V offset on the drive, the HV output (measured directly with a 10x probe) has a 0-(almost)200V triangle wave (for 200V HV in), saturating near 200V and near 0V somewhat before reaching the full range of the triangle
    • The HV mon gives the same answer as measuring the HV output directly, and is reduced 100x compared to the HV output.
    • At 1Hz and above, the rolloff of the low pass still attenuates the drive a bit, and we don't reach the full range.
  • Drove the HV dither pins (1,6) with a 100mV to 10V triangle wave, around 15kHz
    • Even at 10V, the dithering is near the noise of the mon channel, so while I could see a slight peak changing on the FFT near the dither frequency, I couldn't directly observe this on a scope using the mon channel
    • However, measuring the HV directly I do see the dither applied on top of the HV signal. The amplitude of the dither is the same on the HV output as on the dither drive.

[gautam, aaron]

We searched for blips while nominally scanning the OMC length.

We sent a 0.1Hz, 10Vpp triangle wave to the OMC piezo drive diff channels, so the piezo length is seeing a slow triangle wave from 0-200V.

Then, we applied a ~15kHz dither to the OMC length. This dither is added directly onto the HV signal, so the amplitude of the dither at the OMC is the same as the amplitude of the dither into the HV driver.

We monitored the OMC REFL signal (where we saw no blips yesterday) and mixed this with the 15kHz dither signal to get an error signal. Gautam found a pomona box with a low pass filter, so we also low passsed to get rid of some unidentified high frequency noise we were seeing (possibly a ground loop at the function generator? it was present with the box off, but gone with the AC line unplugged). [So we made our own lock-in amplifier.] Photo attached.

We tested the transfer function of the LP, and finding it at 100kHz rather than the advertised 10kHz, we opened the box, removed a resistor to change the 3dB back to 10kHz, and confirmed this by measuring the TF.

We didn't see flashes of error signal in the mixed reflection either, so we suspect that either the PZT is not actuating on the OMC or the alignment is bad. Based on what appears to be the shimmering of far-misaligned fringes on the AS camera, Aaron's suspicion from aligning the cavity with the card, and the lack of flashes, we suspect the alignment. To avoid being stymied by a malfunctioning PZT, we can scan the laser frequency next time rather than the PZT length.

Attachment 1: IMG_4576_copy.jpg
IMG_4576_copy.jpg
  14364   Tue Dec 18 11:42:40 2018 ChubUpdateGeneralAcromag box wired

The Auxiliary DAQ Chassis, or Acromag box, is now wired and ready for testing.  I will be sorting the cables at the vacuum rack to make connection to the box easier.

 

  14363   Mon Dec 17 20:45:40 2018 aaronUpdateOMCAligning the OMC

[gautam, aaron]

We did work in the OMC chamber today to get the OMC aligned. Aaron was in the clean suit while Gautam steered in-air optics. We modified the aux input steering optics and the final two OMC steering optics (between OMMT and OMC), but did not modify any of the AS path optics.

I had already aligned AUX approximately into the AS port from the AP table. With the OMC N door open, we aligned the aux beam first to OM6, then to OMPO, then OM5. OM5 was the last optic in the OMC chamber that we could align to.

From there, Gautam found the aux beam clipping on a few optics on its way to SR4 using the IR viewer. Once we were approximately hitting SR4, we got a return beam in the OMC chamber, which we were able to coalign with the input aux beam.

We had already done the alignment of SR5 into the OMC during the last vent, so we immediately had a refl off of the OMC, which we aligned onto a PD520 from the PSL table (larger aperture than the previous PD, which anyway needed a macroscopic adjustment to catch the refl beam).

Next, we removed the OMC cover, wrapped it in foil, and placed it in the makeshift clean room near the Y end. The screws remain in a foil bucket in the OMC chamber. With the cover off, Aaron moved the OMC input steering mirrors to align the beam in the OMC. We measured ~2.4mW in the OMC refl beam, which means about 240uW is transmitted into the OMC. Aaron thinks the beam overlaps itself after one round trip in the cavity, but that the entire plane may be too low in pitch, so more alignment may be needed here.

With the beam approximately aligned into the OMC, we energized the OMC-L piezo driver with 200V, and applied a ~0.03Hz triangle wave on the OMC diff input (pins 2-7). We monitor the REFL PD, piezo mon, function generator signal, and one of the trans PDs. We noticed that the PZT mon shows the driver saturating before the function generator reaches its full +-10V, which is something to investigate.

We saw what could have been regular dips in the REFL PD signal, but realized that with an unkown level of mode matching, it will be hard to tell whether the light becomes resonant with the DC signal. Gautam has suggested coaligning the aux and PSL beams, then observing the PDH signal from the PSL beam as the OMC sweeps through resonance, while turning aux back on anytime we try to make adjustments to the alignment of the OMC (so I can see the beam in the cavity).

I'll think through the plan in some more detail and we will try to have the OMC locked tomorrow.


gautam:

  1. All references to SR4 etc actually refer to OM4 etc.
  2. For this experiment, we are using the prompt reflection of the AUX beam from the HR surface of the SRM, so as to get maximum light into the cavity.
  3. For 2.4 mW incident on OMC1 (actually we measured the REFL beam on the AS table), we expect ~24 uW inside the cavity, which isn't a lot but still was visible on the card.
  4. After this work, I checked the IMC alignment - it was still easily able to lock to a TEM00 mode, but the transmission was ~half (i.e. ~600cts) of what I am used to it being in low power mode (~1300 cts). I didn't align the cavity to the input beam, as I think in this case, the right thing to do is to align the input beam to the IMC cavity axis.
Attachment 1: IMG_5922.JPG
IMG_5922.JPG
Attachment 2: IMG_5921.JPG
IMG_5921.JPG
  14362   Sat Dec 15 20:04:03 2018 gautamUpdateIOOTT1/TT2 stepping

I'm running a script that moves TT1 and TT2 randomly in some restricted P/Y space to try and find an alignment that gets some light onto the TRY PD. Test started at gpstime 1228967990, should be done in a few hours. The IMC has to remain locked for the duration of this test. I will close the PSL shutter once the test is done. Not sure if the light level transmitted through the ITM, which I estimate to be ~30uW, will be enough to show up on the TRY PD, but worth a shot I figure.

Test was completed and PSL shutter was closed at 1228977122.

  14361   Sat Dec 15 18:29:53 2018 KojiUpdateGeneralChiara new USB 4TB DIsk

The local backup was done at 18:18 after 11h18m of running.

2018-12-15 07:00:01,699 INFO       Updating backup image of /cvs/cds
2018-12-15 18:17:56,378 INFO       Backup rsync job ran successfully, transferred 5717707 files.

 

  14360   Fri Dec 14 22:19:22 2018 KojiUpdateGeneralChiara new USB 4TB DIsk

Edit: It was not 4TB disk but 6TB disk in fact. (We actually ordered 4TB disk...)

I think the problem of the backup disk was the flaky power supply for the external drive.
I swapped the drive to a new HGST 4TB one, but it was neither recognized nor spun up with the external power supply we had. So I decided to put both the new and old drives in the PC chassis to power them up with the internal power supply. I tested the old disk via a USB-SATA cable. However, this disk was not recognized. I noticed that the disk was not HGST 4TB but Seagate 3TB. Is it possible? I thought it was 4TB... Did I miss something?

Once the new 4TB was connected to the USB-SATA, it was very smooth to get it mounted. Now the disk is mounted as /media/40mBackup as before. /etc/fstab was also modified with the new UUID. All the command logs are found here below.

Let's see how the morning backup goes. It would take a while to copy everything on the new disk. So it was actually very nice to set this disk up by Friday midnight.


controls@chiara|~> lsblk

NAME   MAJ:MIN RM   SIZE RO TYPE MOUNTPOINT
sda      8:0    0 465.8G  0 disk 
---- sda1   8:1    0 446.9G  0 part /
---- sda2   8:2    0     1K  0 part 
---- sda5   8:5    0  18.9G  0 part [SWAP]
sdb      8:16   0   1.8T  0 disk 
---- sdb1   8:17   0   1.8T  0 part 
sdc      8:32   0   3.7T  0 disk 
---- sdc1   8:33   0   3.7T  0 part /home/cds
sr0     11:0    1  1024M  0 rom  
sdd      8:64   0   5.5T  0 disk 

controls@chiara|~> sudo mkfs -t ext4 /dev/sdd1

mke2fs 1.42 (29-Nov-2011)
Filesystem label=
OS type: Linux
Block size=4096 (log=2)
Fragment size=4096 (log=2)
Stride=0 blocks, Stripe width=0 blocks
183144448 inodes, 1465130385 blocks
73256519 blocks (5.00%) reserved for the super user
First data block=0
Maximum filesystem blocks=4294967296
44713 block groups
32768 blocks per group, 32768 fragments per group
4096 inodes per group
Superblock backups stored on blocks: 
    32768, 98304, 163840, 229376, 294912, 819200, 884736, 1605632, 2654208, 
    4096000, 7962624, 11239424, 20480000, 23887872, 71663616, 78675968, 
    102400000, 214990848, 512000000, 550731776, 644972544
Allocating group tables: done                            
Writing inode tables: done                            
Creating journal (32768 blocks): done
Writing superblocks and filesystem accounting information: done      

controls@chiara|~> blkid

/dev/sda1: UUID="972db769-4020-4b74-b943-9b868c26043a" TYPE="ext4" 
/dev/sda5: UUID="a3f5d977-72d7-47c9-a059-38633d16413e" TYPE="swap" 
/dev/sdc1: UUID="92dc7073-bf4d-4c58-8052-63129ff5755b" TYPE="ext4" 
/dev/sdd1: UUID="1843f813-872b-44ff-9a4e-38b77976e8dc" TYPE="ext4" 

controls@chiara|~> sudo emacs -nw /etc/fstab
controls@chiara|~> cat /etc/fstab

# /etc/fstab: static file system information.
#
# Use 'blkid -o value -s UUID' to print the universally unique identifier
# for a device; this may be used with UUID= as a more robust way to name
# devices that works even if disks are added and removed. See fstab(5).
#
# <file system> <mount point>   <type>  <options>       <dump>  <pass>
proc            /proc           proc    nodev,noexec,nosuid 0       0
# / was on /dev/sda1 during installation
UUID=972db769-4020-4b74-b943-9b868c26043a /               ext4    errors=remount-ro 0       1
# swap was on /dev/sda5 during installation
UUID=a3f5d977-72d7-47c9-a059-38633d16413e none            swap    sw              0       0
#UUID="90a5c98a-22fb-4685-9c17-77ed07a5e000"    /media/40mBackup       ext4      defaults,relatime,commit=60       0         0
UUID="1843f813-872b-44ff-9a4e-38b77976e8dc"    /media/40mBackup       ext4      defaults,relatime,commit=60       0         0

#fb:/frames      /frames nfs     ro,bg


UUID=92dc7073-bf4d-4c58-8052-63129ff5755b   /home/cds    ext4    defaults,relatime,commit=60    0   0

controls@chiara|~> sudo mount -a
controls@chiara|~> df

Filesystem      1K-blocks       Used  Available Use% Mounted on
/dev/sda1       461229088   10694700  427105320   3% /
udev             15915020         12   15915008   1% /dev
tmpfs             3185412        868    3184544   1% /run
none                 5120          0       5120   0% /run/lock
none             15927044        484   15926560   1% /run/shm
/dev/sdc1      3845709644 1809568856 1840789912  50% /home/cds
/dev/sdd1      5814346836     190408 5521130352   1% /media/40mBackup
  14359   Fri Dec 14 14:25:36 2018 KojiUpdateCDSchiara backup

fsck of chiara backup disk (UUID="90a5c98a-22fb-4685-9c17-77ed07a5e000") was done. But this required many files to be fixed. So the backed-up files are not reliable now.
On the top of that, the disk became not recognized from the machine.

I went to the disk and disconnected the USB and then the power supply, which was/is connected to the UPS.
Then they are reconnected again. This made the disk came back as /media/90a5c98a-22fb-4685-9c17-77ed07a5e000. (*)
After unmounting this disk, I ran "sudo mount -a" to follow the way of mounting as fstab does.
Now I am running the backup script manually so that we can pretend to maintain a snapshot of the day at least.

(*) This is the same situation we found at the recovery from the power shutdown. So my hypothesis is that on Oct 16 at 7 AM during the backup there was a USB failure or disk failure or something which unmounted the disk. This caused some files got damaged. Also this caused the disk mounted as /media/90a5c98a-22fb-4685-9c17-77ed07a5e000. So since then, we did not have the backup.
Update (20:00): The disk connection failed again. I think this disk is no longer reliable.

 

Attachment 1: fsck_log.log
sudo fsck -yV UUID="90a5c98a-22fb-4685-9c17-77ed07a5e000"           [238/276]
[sudo] password for controls:
fsck from util-linux 2.20.1
[/sbin/fsck.ext4 (1) -- /media/40mBackup] fsck.ext4 -y /dev/sde1
e2fsck 1.42 (29-Nov-2011)
/dev/sde1 contains a file system with errors, check forced.
Pass 1: Checking inodes, blocks, and sizes
Error reading block 527433852 (Attempt to read block from filesystem resulted in
 short read) while getting next inode from scan.  Ignore error? yes

... 283 more lines ...
  14358   Fri Dec 14 13:06:12 2018 aaronUpdateOMCAligning the OMC

I replaced the 2'' AUX-AS combining BS with a freshly mounted 2'' HR mirror for 1064. The mirror is labelled 'Y1-2037-45-P', and had a comment on its case: 'V'. I aligned the AUX beam from the new HR mirror into the next iris, so AUX passes through irises both before and after the new optic. Now, AS does not go out to the AS PDs.

I mounted the old BS on the SP table in a random orientation.

I also dumped the beam transmitted through one of the AUX steering mirrors before the new HR mirror.

  14357   Fri Dec 14 13:02:24 2018 aaronUpdateOMCAligning the OMC

I replaced the 2'' AUX-AS combining BS with a 2'' HR mirror for 1064. I aligned the AUX beam from the new HR mirror into the next iris, so AUX passes through irises both before and after the new optic. Now, AS does not go out to the AS PDs.

I mounted the old BS on the SP table in a random orientation.

I also dumped the beam transmitted through one of the AUX steering mirrors before the new HR mirror.

  14356   Thu Dec 13 22:56:28 2018 gautamUpdateCDSFrames

[koji, gautam]

We looked into the /frames situation a bit tonight. Here is a summary:

  1. We have already lost some second trend data since the new FB has been running from ~August 2017.
  2. The minute trend data is still safe from that period, we believe.
  3. The Jetstor has ~2TB of trend data in the /frames/trend folder.
    • This is a combination of "second", "minute_raw" and "minute".
    • It is not clear to us what the distinction is between "minute_raw" and "minute", except that the latter seems to go back farther in time than the former.
    • Even so, the minute trend folder from October 2011 is empty - how did we manage to get the long term trend data?? From the folder volumes, it appears that the oldest available trend data is from ~July 24 2015.

Plan of action:

  1. The wiper script needs to be tweaked a bit to allow more storage for the minute trends (which we presumably want to keep for long term).
  2. We need to clear up some space on FB1 to transfer the old trend data from Jetstor to FB1.
  3. We need to revive the data backup via LDAS. Also summary pages.

BTW - the last chiara (shared drive) backup was October 16 6 am. dmesg showed a bunch of errors, Koji is now running fsck in a tmux session on chiara, let's see if that repairs the errors. We missed the opportunity to swap in the 4TB backup disk, so we will do this at the next opportunity.

  14355   Thu Dec 13 22:36:42 2018 aaronUpdateOMCAligning the OMC

I turned on AUX, and aligned the aux beam to be centered on the first optic the AS beam sees on the AP table. I then turned off the AUX laser.

  14354   Thu Dec 13 22:24:21 2018 aaronUpdateOMCOMC channels

I completed testing of the AI board mentioned above. In addition to the blown fuse, there were two problems:

  • A was a large drop of solder splattered on some of the ch 1 ICs, which is why we couldn't maintain any voltage. I removed the solder.
  • The +12V wire from the power board to the AI board was loose, so I removed and replaced that crimp connection

After this, I tested the TF of all channels. For the most part, I found the expected 3rd order ~7500Hz cheby with notches at ~16kHz and 32kHz. However, some of the channels had shallower or deeper notches. By ~32kHz, I was below the resolution on the spectrum analyzer. Perhaps I just have nonideal settings? I'll attach a few representative examples.

I reinstalled the chassis at 1X2, but haven't connected power.

 

  14353   Thu Dec 13 20:10:08 2018 KojiUpdateGeneralPower Outage recovery

[Gautam, Aaron, Koji]

The PSL interlock system was fixed and now the 40m lab is laser hazard as usual.


- The schematic diagram of the interlock system D1200192
- We have opened the interlock box. Immediately we found that the DC switching supply (OMRON S82K-00712) is not functioning anymore.  (Attachment #1)
- We could not remove the module as the power supply was attached on the DIN rail. We decided to leave the broken supply there (it is still AC powered with no DC output).

- Instead, we brought a DC supply adapter from somewhere and chopped the head so that we can hook it up on the crimping-type quick connects. In Attachment #1, the gray is +12V, and the orange and black lines are GND.

- Upon the inspection, the wires of the "door interlock reset button" fell off and the momentary switch (GRAYHILL 30-05-01-502-03) got broken. So it was replaced with another momentary swicth, which is way smaller than the original unfortunately. (Attachments 2 and 3)

- Once the DC supply adapter was pluged to an AC tap, we heard the sounds of the relays working, and we recovered the laser hazard lamps, PSL door alerm lamps. Also it was confirmed that the PSL innolight is operatable now. 

- BTW, there is the big switch box on the wall close to the PSL enclosure. Some of the green lamps were gone. We found that we have plenty of spare lamps and relays inside of the box. So we replaced the bulbs and know the A.C. lights are functioning. (Attachments 4 & 5)

Attachment 1: OMRON_S82K-00712.JPG
OMRON_S82K-00712.JPG
Attachment 2: reset_button_repaired1.JPG
reset_button_repaired1.JPG
Attachment 3: reset_button_repaired2.JPG
reset_button_repaired2.JPG
Attachment 4: gray_box.JPG
gray_box.JPG
Attachment 5: gray_box2.JPG
gray_box2.JPG
  14352   Thu Dec 13 18:12:47 2018 gautamUpdateIOOND filter on AS camera changed

In order to see the AS beam a bit more clearly in our low-power config, I swapped out the ND=1.0 filter on the AS camera for ND=0.5.

  14351   Thu Dec 13 12:06:35 2018 gautamUpdateGeneralPower Outage recovery

I did a walkaround and checked the status of all the interlock switches I could find based on the SOP and interlock wiring diagram, but the PSL remains interlocked. I don't want to futz around with AC power lines so I will wait for Koji before debugging further. All the "Danger" signs at the VEA entry points aren't on, suggesting to me that the problem lies pretty far upstream in the wiring, possibly at the AC line input? The Red lights around the PSL enclosure, which are supposed to signal if the enclosure doors are not properly closed, also do not turn on, supporting this hypothesis...

I confirmed that there is nothing wrong with the laser itself - i manually shorted the interlock pins on the rear of the controller and the laser turned on fine, but I am not comfortable operating in this hacky way so I have restored the interlock connections until we decide the next course of action...

Quote:
 

The PSL (Edwin) remains in an interlock-triggered state. We are not sure what is causing this, but the laser cannot be powered on until this is resolved.

  14350   Thu Dec 13 10:03:07 2018 ChubUpdateGeneralOMC chamber

Bob, Aaron, and I removed the door from the OMC chamber this morning.  Everything went well.

  14349   Thu Dec 13 01:26:34 2018 gautamUpdateGeneralPower Outage recovery

[koji, gautam]

After several combinations of soft/hard reboots for FB, FEs and expansion chassis, we managed to recover the nominal RTCDS status post power outage. The final reboots were undertaken by the rebootC1LSC.sh script while we went to Hotel Constance. Upon returning, Koji found all the lights to be green. Some remarks:

  1. It seems that we need to first turn on FB
    • Manually start the open-mx and mx services using
      sudo systemctl start open-mx.service 
      sudo systemctl start mx.service
    • Check that the system time returned by gpstime matches the gpstime reported by internet sources.
    • Manually start the daqd processes using
      sudo systemctl start daqd_*
  2. Then fully power cycle (including all front and rear panel power switches/cables) the FEs and the expansion chassis.
    • This seems to be a necessary step for models run on c1sus (as reported by the CDS MEDM screen) to pick up the correct system time (the FE itself seems to pick up the correct time, not sure what's going on here).
    • This was necessary to clear 0x4000 errors.
  3. Power on the expansion chassis.
  4. Power on the FE.
  5. Start the RTCDS models in the usual way
    • For some reason, there is a 1 second mismatch between the gpstime returned on the MEDM screen for a particular CDS model status, and that in the terminal for the host machine.
    • This in itself doesn't seem to cause any timing errors. But see remark about c1sus above in #2.

The PSL (Edwin) remains in an interlock-triggered state. We are not sure what is causing this, but the laser cannot be powered on until this is resolved.

  14348   Wed Dec 12 18:27:07 2018 JonOmnistructureUpgradeAnalog signals, A/D Acromag added to vacuum system

There turned out to be a few analog signals for the vacuum system after all. The TP2/3 foreline pressure gauges were never part of the digital system, but we wanted to add them, as some of the interlock conditions should be predicated on their readings. Each gauge connects to an old Granville-Phillips 375 controller which only has an analog output. Interfacing these signals with the new system required installing an Acromag XT1221 8-channel A/D unit. Taking advantage of the extra channels, I also moved the N2 delivery line pressure transducer to the XT1221, eliminating the need for its separate Omega DPiS32 controller. When the new high-pressure transducers are added to the two N2 tanks, their signals can also be connected.

The XT1221 is mounted on the DIN rail inside the chassis and I have wired a DB-9 feedthrough for each of its three input signals. It is assigned the IP 192.168.114.27 on the vacuum subnet. Testing the channels in situ revealed a subtley in calibrating them to physical units. It was first encountered by Johannes in a series of older posts, but I repeat it here in one place.

An analog-input EPICS channel can be calibrated from raw ADC counts to physical units (e.g., sensor voltage) in two ways:

  1. Via LINR="LINEAR" by setting the engineering-units fields EGUF="[V_max_adc]", EGUL="[V_min_adc]"
  2. Via LINR="NO CONVERSION" by manually setting the gain ASLO="[V/count]" and offset AOFF="[V_offset]"

From the documentation, under the engineering-units method EPICS internally computes:

where EGUF="eng units full scale", EGUL="eng units low", and "full scale A/D counts" is the full range of ADC counts. EPICS automatically infers the range of ADC counts based on the data type returned by the ADC. For a 16-bit ADC like the XT1221, this number is 2^16 = 65,536.

The problem is that, for unknown reasons, the XT1221 rescales its values post-digitization to lie within the range +/-30,000 counts. This corresponds to an actual "full scale A/D counts" = 60,001. If a multiplicative correction factor of 65,536/60,000 is absorbed into the values of EGUF and EGUL, then the first term in the above summation can be corrected. However, the second term (the offset) has no dependence on "full scale A/D counts" and should NOT absorb a correction factor. Thus adjusting the EGUF and EGUL values from, e.g., 10V to 10.92V is only correct when EGUL=0V. Otherwise there is a bias introduced from the offset term also being rescaled.

The generally correct way to handle this correction is to use the manual "NO CONVERSION" method. It constructs calibrated values by simply applying a specified gain and offset to the raw ADC counts:

calibrated val = (measured A/D counts)  x ASLO + AOFF

The gain ASLO="[(V_max_adc - V_min_adc) / 60,001]" and the offset AOFF="0". I have tested this on the three vacuum channels and confirmed it works. Note that if the XT1221 input voltage range is restricted from its widest +/-10V setting, the number of counts is not necessarily 60,001. Page 42 of the manual gives the correct counts for each voltage setting.

  14347   Wed Dec 12 11:53:29 2018 aaronUpdateGeneralPower Outage

At 11:13 am there was a ~2-3 second interruption of all power at the 40m.

I checked that nobody was in any of the lab areas at the time of the outage.

I walked along both arms of the 40m and looked for any indicator lights or unusual activity. I took photos of the power supplies that I encountered, attached. I tried to be somewhat complete, but didn't have a list of things in mind to check, so I may have missed something. 

I noticed an electrical buzzing that seemed to emanate from one of the AC adapters on the vacuum rack. I've attached a photo of which one, the buzzing changes when I touch the case of the adapter. I did not modify anything on the vacuum rack. There is also 

Most of the cds channels are still down. I am going through the wiki for procedures on what to log when the power goes off, and will follow the procedures here to get some useful channels.

Attachment 1: IMG_0033.HEIC
Attachment 2: IMG_1027.HEIC
Attachment 3: IMG_2605.HEIC
  14346   Tue Dec 11 22:50:07 2018 aaronUpdateOMCAligning the OMC

I did the following:

  • Noticed that the OMC rack's power has +-18V, but I had tested the HV driver with +-15V. Maybe fine, something to watch.
  • Checked that nothing but the OMC driver board was in use on the OMC's Sorensen (the QPD whitening board in the OMC rack is not in use, and anyway is labeled +-15V), then turned down the rack voltage from 18 to 15V. Photos attached of AUX_OMC_S Sorensen bank.
  • I hadn't used the alternative dither before. I started by driving the alternative dither with a 10Vpp sine wave at 1-10 Hz. I have both the DC and AC driver mons on a scope.
    • Initially, I only give it 10V at the HV. I don't see much, nor at 30V, while driving with 0-10V sine waves between 0.1-100Hz.
    • In my last log, I hadn't been using the alternative dither.
  • Instead, I switch over to the main piezo drive, which is sent over DB9. Now I see the following on the AC/DC piezo mon channels:
    • Increasing the HV input (increasing in steps from 10-50V) yields 1V at the DC piezo mon for 50V at the HV input.
    • Driving under a few 100s of Hz results in no change to the AC dither mon. Driving <1Hz results in a small (~10% for a 10Vpp drive) at the HV. I didn't take a full transfer function, but it is the thing to do with cds.
    • Changing the drive amplitude changes the AC mon amplitude proportionally
    • At a few kHz, the 10Vpp drive saturates the AC mon.
    • Photos are in order:
      • 1Hz drive, visible on the DC mon channel in green
      • 1kHz drive 10Vpp, visible on the AC mon channel in violet
      • 1kHz drive 5Vpp
      • 5kHz drive 10Vpp, saturates the AC mon channel
Attachment 1: 8323029A-970E-4BEA-833E-77E709300446.jpeg
8323029A-970E-4BEA-833E-77E709300446.jpeg
Attachment 2: C52735BB-0C56-41A1-B731-678CDDCEC921.jpeg
C52735BB-0C56-41A1-B731-678CDDCEC921.jpeg
Attachment 3: 3F8A0B3B-5C7C-4876-B3A7-332F560D554D.jpeg
3F8A0B3B-5C7C-4876-B3A7-332F560D554D.jpeg
Attachment 4: 3DC88B6A-4213-4ABD-A890-7EC317D9EED0.jpeg
3DC88B6A-4213-4ABD-A890-7EC317D9EED0.jpeg
Attachment 5: C0C4F9C0-9574-4A17-9414-B99D6E27025F.jpeg
C0C4F9C0-9574-4A17-9414-B99D6E27025F.jpeg
Attachment 6: A191E1DE-552F-42A5-BED7-246001248BBD.jpeg
A191E1DE-552F-42A5-BED7-246001248BBD.jpeg
  14345   Tue Dec 11 18:20:59 2018 gautamUpdateOptical LeversBS/PRM HeNe is dead

I found that the BS/PRM OL SUM channels were reading close to 0. So I went to the optical table, and found that there was no beam from the HeNe. I tried power-cycling the controller, there was no effect. From the trend data, it looks like there was a slow decay over ~400000 seconds (~ 5 days) and then an abrupt shutoff. This is not ideal, because we would have liked to use the Oplevs as a DC alignment reference during the ventnoI plan to use the AS camera to recover some sort of good Michelson alignment, and then if we want to, we can switch out the HeNe.

*How can I export PDF from NDscope?

Attachment 1: BSOL_dead.png
BSOL_dead.png
  14344   Tue Dec 11 14:33:29 2018 gautamUpdateCDSNDScope

NDscope is now running on pianosa. To be really useful, we need the templates, so I've made /users/Templates/NDScope_templates where these will be stored. Perhaps someone can write a parser to convert dataviewer .xml to something ndscope can understand. To get it installed, I had to run:

sudo yum install ndscope
sudo yum install python34-gpstime
sudo yum install python34-dateutil
sudo yum install python34-requests

 I also changed the pythonpath variable to include the python3.4 site-packages library in .bashrc

Quote:

https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=44971

Let's install Jamie's new Data Viewer

Attachment 1: ndscope.png
ndscope.png
  14343   Tue Dec 11 14:24:18 2018 aaronUpdateOMCAligning the OMC

I set up a function generator to drive OMC-L, and have the two DCPD mons and the OMC REFL PD sent to an oscilloscope. I need to select a cds channel over which to read the REFL signal.

The two DCPD mon channels have very different behaviors on the PD mons at the sat box (see attachment). PD1 has an obvious periodicity, PD2 has less noise overall and looks more white. I don't yet understand this, and whether it is caused by real light, something at the PDs, or something at the sat box.

I've again gone through the operations that will happen with the OMC chamber vented. Here's how it'll go, with some of the open questions that I'm discussing with Gautam or whoever is around the 40m:

  1. Function generator is driving OMC-L. Right now there is one 150V Kepco supply in use, located on the ground just to the right of the OMC rack. I only have plans to power it on while scanning OMC-L, and until the OMC is fully in use the standard practice will be to use this HV with two people in the lab and shut it off after the immediate activities.
    1. To do: Is a second drive necessary for the TT drivers? I don't think it is during this vent, because we will want to align into the OMC with the TTs in a 'neutral' state. I recall that the way the TT drivers are set up, 0V from the dac to the driver is the 'centered' position for all TTs. Unless we want to compensate for some known shift of the chambers during pumpdown, I think this is the TT position we should use while aligning the OMMT into the OMC.
    2. To do: make sure I'm driving the right pins with the function generator. Update: Seems I was driving the right channels, here's the pinout.
  2. We will use the reflection of aux from the SRM to align into the OMC.
    1. Gautam pointed out that I hadn't accounted for the recombination BS for the aux beam being 90-10. This means there's actually something like 300uW of aux onto the OMC, rather than ~3mW. This should still be enough to see on a card, so it is fine.
    2. However, the aux beam is aligned to be colinear with the AS beam when the SRM is misaligned. So the question is whether the wedge on the SRM makes the SRM-reflected aux beam not colinear with the AS beam

 

---------

Talked with Gautam for a good while about the above plan. In trying to figure out why the DCPD sat box appears to have a different TF for the two PDs (seems to be some loose cabling problem at the mons, because wiggling the cables changed this), we determined that the AA chassis also wasn't behaving as expected--driving the expected channels (28-31) with a sine wave yields some signal at the 100Hz driving frequency, but all save ch31 were noisy. We also still saw the 100Hz when the chassis was unplugged. I will continue pursuing this, but in the meantime I'm making an IDE40 to DB37 connector so I can drive the ADC channels directly with the DAC channels I've defined (need to match pinouts for D080303 to D080302). I also will make a new SCSI to DB37 adapter that is more robust than mentioned here. I also need to replace the cable carrying HV to the OMC-L driver, so that it doesn't have a wire-to-wire solder joint.

We moved a razor blade on the AP table so it is no longer blocking the aux beam. We checked the alignment of aux into the AS port. AUX and AS are not colinear anywhere on the AP table, and despite confirming that the main AS beam is still being reflected off of the OMC input mirror, the returning AUX beam does not reach the AP table (and probably is not reaching the OMC). AUX needs to be realigned such that it is colinear with the AS beam. It would be good if in this configuration, the SRM is held close to its position when the interferometer is locked, but the TTs should provide us some (~2.5mrad) actuation. Gautam will do this alignment and I will calculate whether the TTs will be able to compensate for any misalignment of the SRM.

Here is the new plan and minimal things to do for the door opening tomorrow:

  1. Function generator is driving OMC-L.
    1. The PZT mon channel is sent to the oscilloscope.
    2. To do: confirm again that the triangle wave I send in results in the expected triangle wave going to the OMC, using this mon channel.
  2. The OMC REFL signal is being sent to the AP PD. See photo.
    1. Need to align into this PD, but this alignment can be done in air on the AP table.
  3. Monitor the DCPD signals using the TPs from the sat box going to the oscilloscope.
    1. There may be further problems with the sat box, but for the initial alignment into the OMC only the REFL signal is necessary.
    2. Not minimally necessary, but the sat box needs a new case. It has a front, back, and bottom, but no main case, so the board is exposed.
  4. I will move the OMMT-to-OMC steering mirrors while watching the scope for flashes in the REFL signal.

That is the first, minimal sequence of steps, which I plan to complete tomorrow. After aligned into the OMC, the alignment into the DCPDs shouldn't need modification. Barring work needed to align from OMC to DCPDs, I think most other work with the OMC can be done in-air.

  14342   Tue Dec 11 13:48:04 2018 aaronUpdateOMCOMC channels

Koji gave me some tips on testing this board that I wanted to write down, notes probably a bit intermingled with my thoughts. Thanks Koji, also for the DCC and equipment logging!

  • Test the power and AI boards separately with an external supply, ramping the voltage up slowly for each.
  • If it seems the AI board is actually drawing too much current, may need to check its TPs for where a problem might be
    • If it's really extensive may use an IR camera to see what elements are getting too hot
    • Testing in segments will prevent breaking more components
  • Check the regulator that I've replaced
  • The 1 Ohm resistors may have been acting as extra 1A fuses. i need to make sure the resistors I've used to replace them are rated for >1W, if this is the case.
  • Can check the resistance between +-12V and Gnd inputs on the AI board, if there is a short drawing too much current it may show up there.
  • The 7812 may be an appropriate regulator, but the input voltage may need to be somewhat higher than with the low drop regulator that was used before.
  • I want to double check the diagram on the DCC
  14341   Tue Dec 11 13:42:44 2018 KojiUpdateOMCOMC channels

FYI:

D050368 Anti-Imaging Chassis
https://dcc.ligo.org/LIGO-D050368

https://labcit.ligo.caltech.edu/~tetzel/files/equip

D050368 Adl SUS/SEI Anti-Image filter board 
S/N 100-102 Assembled by screaming circuits. Begin testing 4/3/06 
S/N xxx Mohana returned it to the shop. No S/N or traveler. Put in shop inventory 4/24/06 
S/N 103 Rev 01. Returned from Screaming circuits 7/10/06. complete except for C28, C29 
S/N 104-106 Rev 01. Returned from Screaming circuits 7/10/06. complete except for C28, C29 Needs DRV-135’s installed 
S/N 107-111 Rev 02 (32768 Hz) Back from assembly 7/14/06 
S/N 112-113 Rev 03 (65536 Hz) assembled into chassis and waiting for test 1/29/07 
S/N 114 Rev 03 (65536 Hz) assembled and ready for test 020507 


D050512 RBS Interface Chassis Power Supply Board (Just an entry. There is no file)

https://dcc.ligo.org/LIGO-D050512

RBS Interface Chassis Power Board D050512-00

https://labcit.ligo.caltech.edu/~rolf/jayfiles/drawings/D050512-00.pdf
 

 

  14340   Mon Dec 10 19:47:06 2018 aaronUpdateOMCOMC channels

Taking another look at the datasheet, I don't think LM7812 is an appropriate replacement and I think the LM2940CT-12 is supposed to supply 1A, so it's possible the problem actually is on the power board, not on the dewhitening board. The board takes +/- 15V, not +/- 24...

Quote:
 
  1. I identified that the resistors were 1Ohm, and replaced them (though I couldn't find 1Ohm surface mount resistors). I also replaced the voltage regulator in case it was broken. I couldn't find the exact model, so I replaced the LM2940CT-12 with an LM7812, which I think is the newer 12V regulator.

 

  14339   Mon Dec 10 15:53:16 2018 gautamUpdateLSCSwept-sine measurement with DTT

Disclaimer: This is almost certainly some user error on my part.

I've been trying to get this running for a couple of days, but am struggling to understand some behavior I've been seeing with DTT.

Test:

I wanted to measure some transfer functions in the simulated model I set up.

  • To start with, I put a pendulum (f0 = 1Hz, Q=5) TF into one of the filter modules
  • Isolated it from the other interconnections (by turning off the MEDM ON/OFF switches).
  • Set up a DTT swept-sine measurement
    • EXC channel was C1:OMC-TST_AUX_A_EXC
    • Monitored channels were C1:OMC-TST_AUX_A_IN2 and C1:OMC-TST_AUX_A_OUT.
    • Transfer function being measured was C1:OMC-TST_AUX_A_OUT/C1:OMC-TST_AUX_A_IN2.
    • Coherence between the excitation and output were also monitored.
  • Sweep parameters:
    • Measurement band was 0.1 - 900 Hz
    • Logarithmic, downward.
    • Excitation amplitude = 1ct, waveform = "Sine"

Unexplained behavior:

  • The transfer function measurement fails with a "Synchronization error", at ~15 Hz.
    • I don't know what is special about this frequency, but it fails repeatedly at the same point in the measurement.
  • Coherence is not 1 always
    • Why should the coherence deviate from 1 since everything is simulated? I think numerical noise would manifest when the gain of the filter is small (i.e. high frequencies for the pendulum), but the measurement and coherence seem fine down to a few tens of Hz.

To see if this is just a feature in the simulated model, I tried measuring the "plant" filter in the C1:LSC-PRCL filter bank (which is also just a pendulum TF), and run into the same error. I also tried running the DTT template on donatella (Ubuntu12) and pianosa (SL7), and get the same error, so this must be something I'm doing wrong with the way the measurement is being run / setup. I couldn't find any mention of similar problems in the SimPlant elogs I looked through, does anyone have an idea as to what's going on here?

* I can't get the "import" feature of DTT to work - I go through the GUI prompts to import an ASCII txt file exported from FOTON but nothing selectable shows up in DTT once the import dialog closes (which I presume means that the import was successful). Are we using an outdated version of DTT (GDS-2.15.1)?  But Attachment #1 shows the measured part of the pendulum TF, and is consistent with what is expected until the measurement terminates with a synchronization error.


the import problem is fixed - when importing, you have to give names to the two channels that define the TF you're importing (these can be arbitrary since the ASCII file doesn't have any channel name information). once i did that, the import works. you can see that while the measurement ran, the foton TF matches the DTT measured counterpart.


11 Dec 2pm: After discussing with Jamie and Gabriele, I also tried changing the # of points, start frequency etc, but run into the same error (though admittedly I only tried 4 combinations of these, so not exhaustive).

Attachment 1: SimTF.pdf
SimTF.pdf
  14338   Mon Dec 10 12:29:05 2018 aaronUpdateOMCOMC channels

I kept having trouble keeping the power LEDs on the dewhitening board 'on'. I did the following:

1. I noticed that the dewhitening board was drawing a lot of current (>500mA), so I initially thought that the indicators were just turning on until I blew the fuse. I couldn't find the electronics diagrams for this board, so I was using analagous boards' diagrams and wasn't sure how much current to expect to draw. I swapped out for 1A fuses (only for the electronics I was adding to the system).

2. Now the +24V indicator on the dewhitening board wasn't turning on, and the -24V supply was alternatively drawing ~500mA and 0mA in a ~1Hz square wave. Thinking I could be dropping voltage along the path to the board, I swapped out the cables leading to the whitening/dewhitening boards with 16AWG (was 18AWG). This didn't seem to help.

3. Since the whitening board seemed to be consistently powered on, I removed the dewhitening board to see if there was a problem with it. Indeed, I'd burned out the +24V supply electronics--two resisters were broken entirely, and the breadboard near the voltage regulator had been visibly heated.

  1. I identified that the resistors were 1Ohm, and replaced them (though I couldn't find 1Ohm surface mount resistors). I also replaced the voltage regulator in case it was broken. I couldn't find the exact model, so I replaced the LM2940CT-12 with an LM7812, which I think is the newer 12V regulator.
  2. Though this replacement seemed to work when the power board was disconnected from the dewhitening board, connecting to the dewhitening board again resulted in a lot of current draw.
  3. I depowered the board and decided to take a different approach (see)

I noticed that the +/-15V currents are slightly higher than the labels, but didn't notice whether they were already different before I began this work.

I also noticed one pair of wires in the area of 1X1 I was working that wasn't attached to power (or anything). I didn't know what it was for, so I've attached a picture.

Attachment 1: 52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg
52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg
Attachment 2: 545E5512-D003-408B-9F00-55F985966A16.jpeg
545E5512-D003-408B-9F00-55F985966A16.jpeg
Attachment 3: DFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
DFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
  14337   Mon Dec 10 12:11:28 2018 aaronUpdateOMCAligning the OMC

I did some ray tracing and determined that the aux beam will enter the OMC after losing some power in reflection on OMPO (couldn't find this spec on the wiki, I remember something like 90-10 or 50-50) and the SRM (R~0.9), and then transmission through OMPO. This gives us something like 8%-23% of the aux light going to the OMC, depending on the OMPO transmission. This elog tells me the aux power before the recombination BS is ~37mW, ~3.7mW onto SRM, which is consistent with the OMPO being 90-10, and would mean the aux power onto the OMC is ~3mW, plenty for aligning into the OMC.

Since the dewhitening board I'd intended to use isn't working (see elog) , I'm gong to scan the OMC length with a function generator while adjusting the alignment by hand, as was briefly attempted during the last vent.

I couldn't identify a PD on the AP table that was the one I had used during the last vent, I suspect I coopted the very same PD for the arm loss measurements. It is a PDA520, which has a large (100mm^2) area so I've repurposed it again to catch the OMC prompt reflection during the mode scans. I've mounted it approximately where I expect the refl beam to exit the AS chamber.

I brought over the cart that usually lives at 1X1 to help me organize materials near the OMC chamber for opening.

I replaced the banana connectors we'd been using to send HV to the HV driver with soldered wires going to the final locking connector only, so now the 150V is on a safe cable.

I powered up the DCPD sat box and again confirmed that it's working. I sent a 500Hz sine wave through the sat box and confirmed that I can see the signal in the DCPD channels I've defined in cds. I gave the TT and OMC-L PZT channels bad assignments on the ADC (right now, what reads as 'OMC_PZT_MON' is actually the unfiltered output from the sat box, while the DCPD channels are for the filtered outputs of the box), because the way the signals are grouped on the cables I can't attach all of them at once. For this vent, I'll only really need the DCPD outputs, and since I have confirmed that I can read out both of those I'll fix up the HV driver mon channels later.

Attachment 1: B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
  14336   Fri Dec 7 19:42:47 2018 ranaFrogselogcan't upgrade DokuWiki because of PHP / SL7

All of our wikis (except the 40m one which unfortunately got turned into ligo.org mess) use DokuWiki. This now has an auto-upgrade feature through the Admin web interface.

I tried this recently and it fails with this message:

DokuWiki 2018-04-22a "Greebo" is available for download.
 You're currently running DokuWiki Release 2017-02-19e "Frusterick Manners".
! New DokuWiki releases need at least PHP 5.6, but you're running 5.4.16. You should upgrade your PHP version before upgrading!

So we'll have to wait until SL7 (which is what NODUS is running).

I DID do a 'yum upgrade' which updated all the packages. I also installed yum-cron so that the RPM listings get updated daily. But sadly, SL7 only has PHP 5.4.16 (which is a June 2013 release):

> Package php-5.4.16-43.el7_4.1.x86_64 already installed and latest version

  14335   Fri Dec 7 17:04:18 2018 gautamUpdateIOOIMC transmission
  • Power just before PSL shutter on PSL table = 97 +/- 1 mW. Systematic error unknown.
  • Power from IFO REFL on AP table = 40 +/- 1 mW. Systematic error unknown.

Both were measured using the FieldMate power meter. I was hesitant to use the Ophir power meter as there is a label on it that warns against exceeding 100 mW. I can't find anything in the elog/wiki about the measured inesrtion loss / isolation of the input faraday, but this seems like a pretty low amount of light to get back from PRM. The IMC visibility using the MC_REFL DC values is ~87%. Assuming perfect transmission of the 87% of the 97mW that's coupled into the IMC, and assuming a further 5% loss between the Faraday rejected port and the AP table, the Faraday insertion loss would be ~30%. Realistically, the IMC transmission is lower. There is also some part of the light picked off for IPPOS. Judging by the shape of the REFL spot on the camera, it doesn't look clipped to me.

Either way, seems like we are only getting ~half of the 1W we send in on the back of PRM. So maybe it's worth it to investigate the situation in the IOO chamber during this vent.


c1pslc1susaux,c1iool0,caux  crates were keyed. Also, the physical shutter on the PSL NPRO, which was closed last Monday for the Sundance crew filming, was opened and the PMC was locked. PMC remains locked, but there is no light going into the IMC.

  14334   Fri Dec 7 12:51:06 2018 gautamUpdateIMCIMC ringdown fitting

I started putting together some code to implement some ideas we discussed at the Tuesday meeting here. Pipeline isn't setup yet, but i think it's commented okay so if people want to play around with it, the code lives on the 40m gitlab

Model parameters:

  • T+ --- average transmission of MC1 and MC3.
  • T- --- difference in transmission between MC1 and MC3 (this basis is used rather than T1 and T3, because the assumption is that since they were coated in the same coating run, the difference in transmission should be small, even if there is considerable uncertainty in the actual average transmission number.
  • T2 --- MC2 transmission.
  • Lrt --- Round trip loss in the cavity.
  • "sigma" --- a nuisance parameter quantifying the error in the time domain ringdown data.

Simulation:

  • Using these model parameters, calculate some simulated time-domain ringdowns. Optionally, add some noise (assumed Gaussian).
  • Try and back out the true values of the model parameters using emcee - priors were assumed to be uniformly distributed, with a +/- 20% uncertainty around the central value.
  • For a first test, see if there is any improvement in the parameter estimation uncertainty using only transmission ringdown vs both transmission and reflection.

Initial results and conclusions:

  • Attachment #1 - Simulated time series used for this study. The "fit" trace is computed using the median values from the monte-carlo.
  • Attachment #2 - Corner plots showing the distribution of the estimated parameter values, using only transmission ringdown. The "true" values are indicated using the thick blue lines.
  • Attachment #3 - Corner plots showing the distribution of the estimated parameter values, using both transmission and reflection ringdowns.
  • The overall approach seems to work okay. There seems to be only marginal improvement in the uncertainty in estimated parameters using both ringdown signals, at least in the simulation.
  • However, everything seems pretty sensitive to the way the likelihood and priors are coded up - need to explore this a bit more.

Next steps:

  • Add more simulated measurements, see if we can constrain these parameters more tightly. 
  • Use linear error analysis to see if that tells us which measurements we should do, without having to go through the emcee.

There still seems to be some data quality issues with the ringdown data I have, so I don't think we really gain anything from running this analysis on the data I have already collected - but in the future, we can do the ringdown with complete extinguishing of the input light, and repeat the analysis.

As for whether we should clean the IMC mirrors - I'm going to see how much power comes out at the REFL port (with PRM aligned) this afternoon, and compare to the input power. This technique suffers from uncertainty in the Faraday insertion loss, isolation and IMC parameters, but I am hoping we can at least set a bound on what the IMC loss is.

Attachment 1: time_reflAndTrans.pdf
time_reflAndTrans.pdf
Attachment 2: corner_transOnly.pdf
corner_transOnly.pdf
Attachment 3: corner_reflAndTrans.pdf
corner_reflAndTrans.pdf
  14333   Thu Dec 6 17:33:33 2018 JonOmnistructureGeneralN2 line disconnected

I believe I finally have the N2 gauge working correctly. The wiring is unchanged from its original state and the controller has been recalibrated.

After letting the line pressure drop to 0 PSI as indicated by the analog gauge in the drill-press room, I recorded the number of counts read by the Omega controller. Then I pressurized the line to 80 PSI, again indicated by the analog gauge, and recorded the Omega counts again. I entered these two reference points into the controller (automatically determines the gain and offset from these), then confirmed the readings to agree with the anaog gauge as I varied the line pressure.

The two reference points are:

0 PSI  :  13 counts
80 PSI : 972 counts

 

Quote:

[jon, gautam]

In the latest installment in this puzzler: turns out that maybe the trend of the "N2 pressure" channel increasing over the ~3 day timescale it takes a cylinder of N2 to run out is real, and is a feature of the way our two N2 cylinder lines/regulators are setup (for the automatic switching between cylinders when one runs out). In order to test this hypothesis, we'd like to have the line pressure be 0 initially, and then just have 1 cylinder hooked up. 

 

  14332   Thu Dec 6 11:16:28 2018 aaronUpdateOMCOMC channels

I need to hookup +/- 24 V supplies to the OMC whitening/dewhitening boxes that have been added to 1X2.

There are trailing +24V fuse slots, so I will extend that row to leave the same number of slots open.

While removing one +24V wire to add to the daisy chain, I let the wire brush an exposed conductor on the ground side, causing a spark. FSS_PCDRIVE and FSS_FAST are at different levels than before this spark. The 24V sorensens have the same currents as before according to the labels. Gautam advised me to remove the final fuse in the daisy chain before adding additional links.

gautam: we peeled off some outdated labels from the Sorensens in 1X1 such that each unit now has only 1 label visible reflecting the voltage and current. Aaron will post a photo after his work.

  14331   Tue Dec 4 18:24:05 2018 gautamOmnistructureGeneralN2 line disconnected

[jon, gautam]

In the latest installment in this puzzler: turns out that maybe the trend of the "N2 pressure" channel increasing over the ~3 day timescale it takes a cylinder of N2 to run out is real, and is a feature of the way our two N2 cylinder lines/regulators are setup (for the automatic switching between cylinders when one runs out). In order to test this hypothesis, we'd like to have the line pressure be 0 initially, and then just have 1 cylinder hooked up. When we went into the drill-press area, we heard a hiss, turns out that one of the cylinders is leaking (to be fair, this was labelled, but i thought it isn't great to have a higher N2 concentration in an enclosed space). Since we don't need any actuation ability, I valved off the leaky cylinder, and disconnected the other properly functioning one. Attachment #1 shows the current state.

Attachment 1: IMG_7195.JPG
IMG_7195.JPG
  14330   Tue Dec 4 10:38:12 2018 JonOmnistructureUpgradeUpdated Feedthrough List for Vacuum Acromag Chassis

Based on new input from Chub, attached is the revised list of signal cable feedthroughs needed on the vacuum system Acromag crate. I believe this list is now complete.

Attachment 1: acromag_chassis_feedthroughs.pdf
acromag_chassis_feedthroughs.pdf
  14329   Sun Dec 2 19:32:35 2018 ranaUpdateIOOfit times

need to vary start/stop times in fit to test for systematics

  14328   Sun Dec 2 17:26:58 2018 gautamUpdateIMCIMC ringdown fitting

Recently we wondered at the meeting what the IMC round trip loss was. I had done several ringdowns in the winter of 2017, but because the incident light on the cavity wasn't being extinguished completely (the AOM 0th order beam is used), the full Isogaio et. al. analysis could not be applied (there were FSS induced features in the reflection ringdown signal). Nevertheless, I fitted the transmission ringdowns. They looked like clean exponentials, and judging by the reflection signals (see previous elogs in this thread), the first ~20us of data is a clean exponential, so I figured we may get some rough value of the loss by just fitting the transmission data. 

The fitted storage time is 60.8 \pm 2.7 \mu s.However, this number isn't commensurate with the 40m IMC spec of a critically coupled cavity with 2000ppm transmissivity for the input and output couplers.

Attachment #1: Expected storage time for a lossless cavity, with round-trip length ~27m. MC2 is assumed to be perfectly reflecting. The IMC length is known to better than 100 Hz uncertainty because the marconi RF modulation signal is set accordingly. For the 40m spec, I would expect storage times of ~40 usec, but I measure almost 30% longer, at ~60 usec.

Attachment #2: Fits and residuals from the 10 datasets I had collected. This isn't a super informative plot because there are 10 datasets and fits, but to eye, the fits are good, and the diagonal elements of the covariance matrix output by scipy's curve_fit back this up. The function used to fit the t > 0 portions of these signals (because the light was extinguished at t=0 by actuating on the AOM) is \text{Transmission} = Ae^{-\frac{2t}{\tau_{\mathrm{storage}}}}, where A and tau are the fitted parameters. In the residuals, the same artefacts visible in the reflection signal are seen.

Attachment #3: Scatter plot of the data. Width of circles are proportional to fit error on individual measurements (i just scaled the marker size arbitrarily to be able to visually see the difference in uncertainty, the width doesn't exactly indicate the error), while the dahsed lines are the global mean and +/- 1 sigma levels.

Attachment #4: Cavity pole measurement. Using this, I get an estimate of the loss that is a much more believable 300 \pm 20\, \mathrm{ppm}.

Attachment 1: tauTheoretical.pdf
tauTheoretical.pdf
Attachment 2: ringdownFit.pdf
ringdownFit.pdf
Attachment 3: ringdownScatter.pdf
ringdownScatter.pdf
Attachment 4: cavPole.pdf
cavPole.pdf
  14327   Sun Dec 2 16:08:44 2018 JonOmnistructureUpgradeFeedthroughs for Vacuum Acromag Chassis

Below is an inventory of the signal feedthroughs that need to be installed on the vacuum Acromag crate this week.

Type Qty Connects to # Chs Signals
DB-37 female 1 Main AC relay box 18 Valve/roughing pump control
DB-9 female 5** Satellite AC relay boxes 3-4/box Valve control
DB-25 male 1 Turbo pump 1 controller 5 Pump status readbacks
DB-9 male 30 Valve position indicators 2/valve Valve position readbacks
DB-9 male 3 Roughing pump controllers 1/pump Pump status readbacks
DB-9 male 1 Cryo pump controller 2 Pump status readbacks

**The original documentation lists five satellite boxes (one for each test mass chamber and one for the beamsplitter chamber), but Chub reports not all of them are in use. We may remove the ones not used.

  14326   Fri Nov 30 19:37:47 2018 gautamUpdateLSCLSC feedforward block diagram

I wanted to set up an RTCDS model to understand this problem better. Attachment #1 is the simulink diagram of the signal flow. The idea will be to put in the appropriate filter shapes into the various filter blocks denoting the DARM and auxiliary DoF plants, controllers and actuators, and then use awggui / diaggui to inject some noises and see if in this idealized model I can achieve good subtraction. Then we can build up to applying a time varying cross coupling between DARM and the vertex DoF, and see how good the adaptive FF works. Still need to setup some MEDM screens to make working with the test system easier.

I figured c1omc would be the least invasive model to set this upon without risking losing any of our IR/green alignment references. Compile and install went smooth, see Attachment #2. The c1omc model was clocking 4us before, now it's using 7us.

Attachment #3 shows the top level of the OMC model, while Attachment #4 shows the MEDM screen.

* Note to self: when closing a loop inside the realtime model, there has to be a delay block somewhere in the loop, else a compilation error is thrown.

Attachment 1: LSC_FF_tester.png
LSC_FF_tester.png
Attachment 2: Screenshot_from_2018-11-30_19-41-07.png
Screenshot_from_2018-11-30_19-41-07.png
Attachment 3: Screenshot_from_2018-12-10_15-31-23.png
Screenshot_from_2018-12-10_15-31-23.png
Attachment 4: SimLSC.png
SimLSC.png
ELOG V3.1.3-